OpenStudy (anonymous):

prove the ff. identities 1. (sin 1/2x-cos1/2x)^2= 1-sin x 2. csc x/sec x= 1 + cot x/1 + tan x

hartnn (hartnn):

u know the formula for (a+b)^2 ?

OpenStudy (anonymous):

nope, i forgot all of them when i was in high school

hartnn (hartnn):

for 1. i will use these three formulas, write them down. \(\large (a+b)^2=a^2+2ab+b^2------>(1)\\\large2sin xcos x=sin2x----->(2)\\\large sin^2x+cos^2x=1------>(3)\)

OpenStudy (anonymous):

ok

hartnn (hartnn):

now take left side: \((sin (x/2)-cos(x/2))^2\) expand it accordind to (1) and tell me what u get.

OpenStudy (anonymous):

if you square sin x/2, is the answer sin^2 x/2?

hartnn (hartnn):

yes.

OpenStudy (anonymous):

then the middle term will be -2sin cos x/2?

hartnn (hartnn):

-2sin(x/2)cos(x/2)

OpenStudy (anonymous):

so sin^2 x/2- 2 sin(x/2)cos(x/2)- cos^2 x/2

hartnn (hartnn):

careful with brackets and signs. \(sin^2 (x/2)- 2 sin(x/2)cos(x/2)+ cos^2 (x/2)\)

OpenStudy (anonymous):

how did you use that italicized thing?

hartnn (hartnn):

now use (3) formula, what will be \(sin^2 (x/2)+ cos^2 (x/2)=??\)

OpenStudy (anonymous):

wait.

hartnn (hartnn):

`write that in \(...........\)` like `\(sin^2 x\)` u get \(sin^2x\)

OpenStudy (anonymous):

that would be equal to 1?

hartnn (hartnn):

yes! thats one :) so u have now : \(1−2sin(x/2)cos(x/2)\) ok?

OpenStudy (anonymous):

so 1-2sin(x/2)cos(x/2) is equal to 1-sin2(x/2)?

hartnn (hartnn):

YES! from the (2) formula. since 2sinx cos x = sin 2x 2sin(x/2)cos(x/2) = sin 2(x/2) =sin x

OpenStudy (anonymous):

thats the final answer?

OpenStudy (anonymous):

what will happen to 1-sin x?

hartnn (hartnn):

u just proved that (sin 1/2x-cos1/2x)^2= 1-sin x should i put all steps together?

OpenStudy (anonymous):

yes

hartnn (hartnn):

\( (sin(x/2)-cos(x/2))^2=sin^2 (x/2)- 2 sin(x/2)cos(x/2)+ cos^2 (x/2)\\=1-2 sin(x/2)cos(x/2)=1-sin x\)

hartnn (hartnn):

u have list of trignometric formulas?

OpenStudy (anonymous):

i have them, but they are a lot. what specific identities? like fundamental identities?

hartnn (hartnn):

i'll list them: \(\huge sin x =1/csc x \implies csc x=1/sin x \\\huge cos x = 1/sec x \implies secx=1/cos x \\\huge tan x=1/cot x \implies cot x =1/tanx \\\huge tan x=sin x/cos x \\\huge cot x=cos x/sinx \\ \huge sin^2x+cos^2x=1 \\\huge sec^2x=1+tan^2x\\\huge cosec^2x=1+cot^2x\)

hartnn (hartnn):

here u will not need last 3,but they are VERY IMPORTANT

OpenStudy (anonymous):

what do u mean?

hartnn (hartnn):

all these are important, for 2. u'll need few from first 5.

hartnn (hartnn):

so lets take right side first, can u see cot x and tan x ? what will u write those as ?

OpenStudy (anonymous):

tan x= 1/cot x -------cot x=1/tan x

hartnn (hartnn):

that will be of no use, try cosx/sinx and sinx/cos x

hartnn (hartnn):

\(\huge \frac{1+cotx}{1+tan x}=\frac{1+\frac{cos x}{sin x}}{1+\frac{sinx}{cosx}}=?\)

OpenStudy (anonymous):

1+ cosx/sin x=1 + cot x

OpenStudy (anonymous):

and 1 +tan x= 1+ sinx/cos x

hartnn (hartnn):

??

OpenStudy (anonymous):

so 1+cot x/1+tan x= 1+ cot x/1 +tan x

hartnn (hartnn):

lol!

OpenStudy (anonymous):

what is wrong?

hartnn (hartnn):

i wrote cot x as cos x/sin x and u again wrote cos x/sin x as cot x!!! so whats the use of my writing cot x as cos x/sin x ??

OpenStudy (anonymous):

sorry, i just realized

OpenStudy (anonymous):

csc x/sec x= 1 + cot x/ 1+ tan x

hartnn (hartnn):

ok, i will write next step, see whether u get it: \(\huge \frac{1+\frac{cos x}{sin x}}{1+\frac{sinx}{cosx}}=\frac{(sin x +cos x)/sinx}{(sin x+cos x)/cosx}\)

OpenStudy (anonymous):

are we working for the two sides?

hartnn (hartnn):

i started from 1+cotx/1+tanx and i will reach, csc x/sec x

hartnn (hartnn):

now can u see that (sin x +cos x) gets cancelled from numerator and denominator....

OpenStudy (anonymous):

yes

hartnn (hartnn):

so what remains ?

OpenStudy (anonymous):

cos x/sin x

hartnn (hartnn):

absolutely correct, but i will keep it in the form of : \(\huge \frac{1/sin x}{1/cos x}=?\) now see the formula list and tell me what is 1/sin x and 1/cos x ??

OpenStudy (anonymous):

1/sinx= csc x and 1/cos x= sec x

OpenStudy (anonymous):

can u write the whole steps?

hartnn (hartnn):

and u are done!! see all mt comments with large font....u will get it.

hartnn (hartnn):

*my

OpenStudy (anonymous):

can u write it?

hartnn (hartnn):

ok :) \(\huge \frac{1+\frac{cos x}{sin x}}{1+\frac{sinx}{cosx}}=\frac{(sin x +cos x)/sinx}{(sin x+cos x)/cosx}=\huge \frac{1/sin x}{1/cos x}=\frac{cscx}{secx}\)

OpenStudy (anonymous):

thanks :)

hartnn (hartnn):

welcome :)